Kabbed:

Angie is working on solving the exponential equation 23x = 6; however, she is not quite sure where to start. Using complete sentences, describe to Angie how to solve this equation.

Kabbed:

is it called moving down the exponent

Kabbed:

putting the x

Hero:

Is that the original problem? You should post the original problem.

Kabbed:

i dont need the answer but i changed the question

Kabbed:

Since were following the laws of log we put a x but what is it called since I have to explain is it called moving the exponent down or something

Hero:

Okay, great. So the exponential equation is \(23^x = 6\) And you started off correctly by applying the log to both sides of the equation to get \(\log(23^x) = \log(6)\) The next step would be to apply what is called the \(\textbf{Power Property Of Logarithms}\) which states $$\log(a^b)= b \log(a)$$

Hero:

And that is what you would say to describe how you get from the previous step where you applied the log to both sides to this point: \(x \log(23) = \log(6)\)

Hero:

Any questions?

Kabbed:

it states log(ab)=blog(a) but i dont see it state any x

Kabbed:

oh wait I see it brings the exponent down

Hero:

The \(b\) IS the \(x\)

Kabbed:

Wow it makes a lot more sense now thank you so much for helping me bye

Hero:

You're welcome